OpenStudy (loser66):

\(F=(\dfrac{x}{r}, \dfrac{y}{r},\dfrac{z}{r})\) where \(r=\sqrt{x^2+y^2+z^2}\) Find div F, please, help

OpenStudy (loser66):

@misty1212

OpenStudy (baru):

if F=<P,Q,R> \[Div(F)=P_x +Q_y +R_z\]

OpenStudy (baru):

you might simplify calculations using the chain rule for partial derivatives P=P(x,r) \[\frac{\partial P}{\partial x} =(\frac{\partial P}{\partial x})_r+\frac{\partial P}{\partial r}\frac{\partial r}{\partial x}\]

OpenStudy (hulahoop):

you have \(<\partial_x, \partial_y, \partial_z> \bullet <\dfrac{x}{r}, \dfrac{y}{r}, \dfrac{z}{r}>\) looking purely at the \(\partial_x \left( \dfrac{x}{r}\right)\) bit, as the symmetry does the rest, you can go at it from quotient rule so \(\left. \partial_x \left( \dfrac{x}{r}\right) \right|_{y,z = const}= \dfrac{\partial_x(x)r - x \partial_x(r)}{r^2} = \dfrac{r - x. \dfrac{x}{r}}{r^2}\) \(=\dfrac{r^2 - x^2}{r^3}\) and where \(\partial_x (r) = \left. \partial_x (\sqrt{x^2 + y^2 + z^2}) \right|_{y,z = const} = \dfrac{x}{r}\) repeat for y and z, and add it all up.... i am sure you could also switch to spherical div, and maybe do this more efficiently.

OpenStudy (loser66):

Thanks for that. since F is symmetric so we have the same for y, z hence the result is \(\dfrac{1}{r^3} \) but it is not, the result is 2/r. I don't know how

OpenStudy (hulahoop):

from \[\left. \partial_x \left( \dfrac{x}{r}\right) \right|_{y,z = const}= \dfrac{\partial_x(x)r - x \partial_x(r)}{r^2} = \dfrac{r - x. \dfrac{x}{r}}{r^2}\] you get \[ \partial_x \left( \dfrac{x}{r}\right) = \dfrac{r^2 - x^2}{r^3}\] so applying the symmetry: \( \nabla \bullet <\dfrac{x}{r}, \dfrac{y}{r}, \dfrac{z}{r}> = \dfrac{r^2 - x^2}{r^3} + \dfrac{r^2 - y^2}{r^3} + \dfrac{r^2 - z^2}{r^3} = \dfrac{3r^2 - r^2}{r^3}\) \(= \dfrac{2r^2 }{r^3} = \dfrac{2 }{r}\)

OpenStudy (loser66):

Thank you so much.